Question #1: If a preparation contains 5grams of a drug in 500 mL. how many grams are contained in each tablespoon?: Question #1: If a preparation contains 5grams of a drug in 500 mL. how many grams are contained in each tablespoon?
Answer by nerdybill(1122)

The key to this problem is to determine the volume of a single tablespoon (in metric units).
.
Reference:
http://www.joyofbaking.com/USMetricVolumeEquiv.html
.
1 tablespoon = 15 ml
.
Your proportion is then:
Let x = grams of drug in a teaspoon
5/500 = x/15
x = (5*15)/500
x = 75/500
x = 0.15 grams (or, 150 milligrams)